Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, November 4, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 4, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/04/98 (Wednesday)

Segment 1 (Emidio to ANF): Report for 11/4/98: De-watering of the 8 mile segment of pipe from EMP's 33.34 to 25.48 continued. On Tejon Bench, excavation of trench began again at EMP 18.40 working south. Progress through here will be slow due to limited work space. Backfill is completed to ~18.39. Shading is taking place from EMP's 17.45 to 17.42. Excavation of trench continued from ~EMP 17.14 working north. In Grapevine Canyon the road has been sawed from EMP 16.43 to 15.73. Excavation of the trench began at 16.43 working north. At Grapevine Pump Station electrical work on the pumps continued. Pipe fitting in the pipe vault and foundation work on the sub-station progressed. The bore at the California Aqueduct continued (EMP 12.28). Shoring was being welded at the Stream Crossing #4 bore (EMP 9.03). On Road B, pipe has been welded and x-rayed from EMP 6.38 to 6.12. The coating crew worked to EMP 6.53. On Hwy. 166 excavation continued from EMP .65 working west. Drilling of the bore at EMP 1.18 progressed. At Emidio Pump station electrical work and pipe fitting continued. Electrical wiring took place in the office building.

Segment 2 (ANF to Sylmar): 2A (Copco Rd to Templin Hwy): The testing cutouts at Liebre, EMP 42.27, and Forest Inn site, 47.35, have been tied-in and backfilled. Equipment was hauled away from the Liebre site today. 2C (Templin Hwy to Castaic): Work continued on the block wall on the Whitaker Reduction Station access road. The electrical sub-contractor finished work at the Station today. Straw was spread over previously drilled seed between stream crossing #46, EMP 53.00, and I-5 bore, 53.45. The steepest slopes in this section will be revegetated with curlex beginning tomorrow. Revegetation with curlex also took place at EMP 55.06. Welding in the trench moved north to EMP 55.51, x-ray to 55.58, and coating to 55.84. Shading reached 56.03. 2B (Castaic to Sylmar): Fusing of F/O conduit occurred on The Old Road at Hwy 126. Crews began the set-up of the directional drill at Feedmill Rd. Welding, coating, and insertion of pipe into the casing took place at Calgrove Blvd, EMP 69.08. The roadway was ground in order to set steel plates flush over open trench north of Sierra Hwy at EMP 72.55. A pull box for F/O conduit was set at EMP 73.0. Backfill with base rock occurred between EMP 73.1 and 73.19. 300 feet of asphalt was laid at the same location. Pipe was bent, slung, welded, and x-rayed under the I-5 bridge north of Sunshine Cyn. at EMP 73.25. Trench was excavated and dewatered next to the Mobile oil line at Sunshine Cyn, EMP 73.34.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 4 November 1998: 3A: Backfill crews moved from the Downey Bridge to begin work on the backfill of Verdugo Wash. Other work consisted of fiber optic conduit being pulled between Verdugo and Riverdale in Burbank. Ditch work on the north side of the Hollywood Way Bridge continued. Bore pit excavation began on the south side of Tuxford Street. NIGHT work was performed on the San Fernando Bridge crossing, including welding, x-ray, and coating. 3B: Razor wire continued to be installed on the Downey Bridge. Crews began marking the location of the line, in Midway Yard, for markers to be installed. 3C: The bore pit at Hubbard Avenue was completely excavated. The bore crew set up and was waiting for the OSHA inspection. Backfill worked northward to near MP 74.71. Ditch crews neared completion of their work in San Fernando Road, while the crew on the MTA right of way completed their work up to the bore site. Frame work continued on the California Aqueduct crossing.

Segment 4B (El Segundo to Lynwood) 11-4-98: Backfilling activities continued at 108 Th St. ESEMP 1.16, Santa Ana ESEMP 0.04, Alameda and Fernwood AV ESEMP 0.02, and the crew at Western ESEMP 5.56 was backfilling the bore pit. Construction activities continued at El Segundo station ESEMP 11.00. This crew was working on getting the soil out of the casing. The crew at Lynwood station EMP 0.0. was installing pips today. All the backfilling at Dominguez Channel has been completed

Segment 5 (Lynwood to Long Beach) Report for 4 November 1998: Pilot hole for directional drill at Highway 91/Compton Creek is almost through. Frac-outs (upwellings of boring mud) slowed progress toward the end of the day. Crew at Wilmington station is preparing for bore into refinery.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page